# Quantum Algorithms for the Pathwise Lasso

### v2

@@ 4.4 Gaussian random design matrix
In this section, we shall bound these quantities for the specific case when the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ is a random matrix from the $\mathbf{\Sigma}$-Gaussian ensemble. According to Fact 22, in this situation the Lasso solution is unique almost surely, thus $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ and $\mathbf{X}_{\mathcal{A}}^{+}=(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top}$. As we shall prove below, with high probability, $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$ is $O(\sqrt{\log(d)/n)}$ and the mutual incoherence is at most $1/2$. The mutual overlap between $\mathbf{y}$ and $\mathbf{X}$, on the other hand, is $\Omega(\|\mathbf{y}\|_{2}/(\|\mathbf{y}\|_{1}\sqrt{\log{d}}))$ with high probability, which equals $\Omega(1/\sqrt{\log{d}})$ when $\mathbf{y}$ is sparse. These bounds yield quantum and classical complexities $\widetilde{O}(\sqrt{d}/\epsilon)$ and $\widetilde{O}(d/\epsilon^{2})$ per iteration for Algorithm 3 and Algorithm 4, respectively, as stated in Result 4.
@@
By a Chernoff bound (Fact 6) plus a union bound, we have that $\mathbb{P}[\|\mathbf{Z}\|_{\max}\leq\sqrt{\ln(4nd/\delta)}]\geq 1-\delta/2$, which implies that $\mathbb{P}[\|\mathbf{X}\|_{\max}\leq\sqrt{\ln(4nd/\delta)}\|\sqrt{\mathbf{\Sigma}}\|_{1}]\geq 1-\delta/2$.
We now focus on $\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$. The rows of $\mathbf{X}_{\mathcal{A}}$ are i.i.d. sampled from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{\Sigma}_{\mathcal{A}|\mathcal{A}^{c}})$. According to [114, Theorem 6.1],
@@
We now move on to $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}$. The next lemma tells us that, if $|\mathcal{A}|=O(\sigma_{\min}(\mathbf{\Sigma})n/\log{d})$, where $\sigma_{\min}(\mathbf{\Sigma})$ is the minimum eigenvalue of $\mathbf{\Sigma}$, then the mutual incoherence is a constant bounded away from $1$ with high probability. A similar result can be found in [114, Exercise 7.19].
@@
Consider the quantity $|\mathbf{X}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|$ for fixed $j\in\mathcal{A}^{c}$. The zero-mean Gaussian vector $\mathbf{X}_{j}$ can be decomposed into a linear prediction based on $\mathbf{X}_{\mathcal{A}}$ plus a prediction error as (cf. [113, Section V.A] and [114, Equation (11.26) & Exercise 11.3])
@@
Since $\mathbf{W}_{j}$ is independent of $\mathbf{X}_{\mathcal{A}}$, $\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}$ is a sub-Gaussian random variable with parameter (Fact 7)
@@
where we used that $[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj}\leq 1$ since $\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}\preceq\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}^{c}}$ and the diagonal entries of $\mathbf{\Sigma}$ are at most $1$. We now relate $\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{2}$ with $\sigma_{\min}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})\geq\sigma_{\min}(\mathbf{\Sigma})$. According to [113, Lemma 9],
@@
$\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}\succeq C_{\mathbf{\Sigma}}\mathbf{I}\quad\text{for}\leavevmode\nobreak\ C_{\mathbf{\Sigma}}>0$
@@
Let us start by bounding $\|\mathbf{X}\|_{\max}$. Similarly to Lemma 31, $\|\mathbf{X}\|_{\max}\leq\|\sqrt{\mathbf{\Sigma}}\|_{1}\|\mathbf{Z}\|_{\max}$, where $\mathbf{Z}$ is drawn from the $\mathbf{I}$-Gaussian ensemble. By a Chernoff bound (Fact 6) plus a union bound, $\mathbb{P}[\|\mathbf{Z}\|_{\max}\leq\sqrt{\ln(6nd/\delta)}]\geq 1-\delta/3$, which implies that $\mathbb{P}[\|\mathbf{X}\|_{\max}\leq\sqrt{\ln(6nd/\delta)}\|\sqrt{\mathbf{\Sigma}}\|_{1}]\geq 1-\delta/3$.
@@
since $[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj}\geq C_{\mathbf{\Sigma}}$. Furthermore, since $\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+}$ is a random projection, i.e., a projection onto a random $(n-|\mathcal{A}|)$-dimensional subspace uniformly distributed in the Grassmann manifold $G_{n,n-|\mathcal{A}|}$, then Lévy’s lemma (Fact 9) yields [^4] The function $\mathbf{u}\in\mathbb{R}^{n}\mapsto\|\mathbf{P}\mathbf{u}\|_{2}$ is $1$-Lipschitz for any orthogonal projector $\mathbf{P}\in\mathbb{R}^{n\times n}$. (see also [110, Lemma 5.3.2])

### v3

@@ 4.4 Gaussian random design matrix
In this section, we shall bound these quantities for the specific case when the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ is a random matrix from the $\mathbf{\Sigma}$-Gaussian ensemble. According to Fact 19, in this situation the Lasso solution is unique almost surely, thus $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ and $\mathbf{X}_{\mathcal{A}}^{+}=(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top}$. As we shall prove below, with high probability, $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$ is $O(\sqrt{\log(d)/n)}$ and the mutual incoherence is at most $1/2$. The mutual overlap between $\mathbf{y}$ and $\mathbf{X}$, on the other hand, is $\Omega(\|\mathbf{y}\|_{2}/(\|\mathbf{y}\|_{1}\sqrt{\log{d}}))$ with high probability, which equals $\Omega(1/\sqrt{\log{d}})$ when $\mathbf{y}$ is sparse. These bounds yield quantum and classical complexities $\widetilde{O}(\sqrt{d}/\epsilon)$ and $\widetilde{O}(d/\epsilon^{2})$ per iteration for Algorithms 3 and 4, respectively, as stated in Result 3.
@@
By a Chernoff bound (Fact 3) plus a union bound, we have that $\mathbb{P}[\|\mathbf{Z}\|_{\max}\leq\sqrt{\ln(4nd/\delta)}]\geq 1-\delta/2$, which implies that $\mathbb{P}[\|\mathbf{X}\|_{\max}\leq\sqrt{\ln(4nd/\delta)}\|\sqrt{\mathbf{\Sigma}}\|_{1}]\geq 1-\delta/2$.
We now focus on $\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$. The rows of $\mathbf{X}_{\mathcal{A}}$ are i.i.d. sampled from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{\Sigma}_{\mathcal{A}|\mathcal{A}^{c}})$. According to [116, Theorem 6.1],
@@
We now move on to $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}$. The next lemma tells us that, if $|\mathcal{A}|=O(\sigma_{\min}(\mathbf{\Sigma})n/\log{d})$, where $\sigma_{\min}(\mathbf{\Sigma})$ is the minimum eigenvalue of $\mathbf{\Sigma}$, then the mutual incoherence is a constant bounded away from $1$ with high probability. A similar result can be found in [116, Exercise 7.19].
@@
Consider the quantity $|\mathbf{X}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|$ for fixed $j\in\mathcal{A}^{c}$. The zero-mean Gaussian vector $\mathbf{X}_{j}$ can be decomposed into a linear prediction based on $\mathbf{X}_{\mathcal{A}}$ plus a prediction error as (cf. [115, Section V.A] and [116, Equation (11.26) & Exercise 11.3])
@@
Since $\mathbf{W}_{j}$ is independent of $\mathbf{X}_{\mathcal{A}}$, $\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}$ is a sub-Gaussian random variable with parameter (Fact 4)
@@
where we used that $[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj}\leq 1$ since $\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}\preceq\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}^{c}}$ and the diagonal entries of $\mathbf{\Sigma}$ are at most $1$. We now relate $\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{2}$ with $\sigma_{\min}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})\geq\sigma_{\min}(\mathbf{\Sigma})$. According to [115, Lemma 9],
@@
$\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}\succeq C_{\mathbf{\Sigma}}\mathbf{I}\quad\text{for}\quad C_{\mathbf{\Sigma}}>0,$
@@
Let us start by bounding $\|\mathbf{X}\|_{\max}$. Similarly to Lemma 29, $\|\mathbf{X}\|_{\max}\leq\|\sqrt{\mathbf{\Sigma}}\|_{1}\|\mathbf{Z}\|_{\max}$, where $\mathbf{Z}$ is drawn from the $\mathbf{I}$-Gaussian ensemble. By a Chernoff bound (Fact 3) plus a union bound, $\mathbb{P}[\|\mathbf{Z}\|_{\max}\leq\sqrt{\ln(6nd/\delta)}]\geq 1-\delta/3$, which implies that $\mathbb{P}[\|\mathbf{X}\|_{\max}\leq\sqrt{\ln(6nd/\delta)}\|\sqrt{\mathbf{\Sigma}}\|_{1}]\geq 1-\delta/3$.
@@
since $[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj}\geq C_{\mathbf{\Sigma}}$. Furthermore, since $\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+}$ is a random projection, i.e., a projection onto a random $(n-|\mathcal{A}|)$-dimensional subspace uniformly distributed in the Grassmann manifold $G_{n,n-|\mathcal{A}|}$, then Lévy’s lemma (Fact 6) yields [^4] The function $\mathbf{u}\in\mathbb{R}^{n}\mapsto\|\mathbf{P}\mathbf{u}\|_{2}$ is $1$-Lipschitz for any orthogonal projector $\mathbf{P}\in\mathbb{R}^{n\times n}$. (see also [112, Lemma 5.3.2])